OpenStudy (anonymous):

Can someone please tell me if my work is right or wrong, my teacher says I need to re do some problems but I am not sure what it wrong.

OpenStudy (anonymous):

1. 2x^4(4x^2+ 3x + 1) = 8x^6+6x^5+2 2x^4 x 4x^2= 8^6 2x^4 x 3x^1= 6^5 2x^4 x 1= 2 2. (4x^1 – 3)(2x^2– 7x^1 + 1) = 8x^3-34x^2+21x+4-3 4x^1(2x^2– 7x^1 + 1)= 8x^3-28^2+4 -3(2x^2– 7x^1 + 1) = -6x^2+21x-3 3. (x2+ 4x – 3)(2x2+ x + 6) (2x^4 + x^3 +6x^2) + (8x^3 + 4x^2 + 24x) + (-6x^2 - 3x -18)  2x^4 + 9x^3 +4x^2 +21x - 18 4. Write a simplified polynomial expression to represent the area of the rectangle below. (x + 5) (2x - 4) p(x) = 2x² - 4x + 10x - 20 p(x) = 2x² + 6x - 20 5. Write a simplified polynomial expression to represent the area of the square tile, shown below. 4 (2x-4)^2+(x+5)^2

OpenStudy (johnweldon1993):

For number 1 \[\large 2x^4 \times 1 \cancel{=} 2\]

OpenStudy (anonymous):

I don't understand?

OpenStudy (johnweldon1993):

\[\large 2x^4(4x^2+ 3x + 1)\] \[\large 2x^4 \times 4x^2 = 8x^6\] \[\large 2x^4 \times 3x = 6x^5\] \[\large \color{red}{2x^4 \times 1 = 2x^4}\]

OpenStudy (anonymous):

So for Number 1 I forgot to put the x with 2, okay got I understand that one now thank you

OpenStudy (johnweldon1993):

And for number 2 now \[\large (4x^1 – 3)(2x^2– 7x^1 + 1)\] \[\large 4x \times 2x^2 = 8x^3\] \[\large 4x \times -7x = -28x^2\] \[\large 4x \times 1 = 4x\] \[\large -3 \times 2x^2 = -6x^2\] \[\large -3 \times -7x = 21x\] \[\large -3 \times 1 =-3\] so we have \[\large 8x^3 - 28x^2 - 6x^2 + 4x + 21x - 3\] \[\large \color{red}{8x^3 - 34x^2 + 25x - 3}\]

OpenStudy (anonymous):

It makes more sense now, I think I got confused when doing the second part and i think I added wrong

OpenStudy (anonymous):

I had 6x^2 when there shouldn't have been 6x^2 at all